# A Survey on Speech Large Language Models

### v2

@@ 3.1 Speech LLM Input-Output Modes
    The ST2T mode, the most widely used in modern speech recognition, incorporates both audio and text inputs. Typically, the text input serves as an instruction or task prompt. The model processes both modalities, fusing the audio and text information to produce a final text output. This approach enables multi-task capabilities and allows models to fully leverage the strengths of large language models, handling a wider variety of tasks than the S2T mode.
@@
    In this more advanced mode, a vocoder is added to the output, enabling the model to generate both text and audio outputs. This expands the model’s capabilities beyond basic speech recognition, supporting tasks such as Text-to-Speech (TTS), speech translation, and voice conversion. The ST2ST mode offers greater versatility than the other two modes, making it applicable across a broader range of speech and language tasks.
3.2 Model structure
Fig. 3: Overview of key stages of the Speech LLM Workflow
@@ 3.2 Model structure
Up to now, various Speech LLM architectures have been developed, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference. As illustrated in Figure 3, despite different input and output modalities, the major stages that a model went through is consistent.
Speech LLMs process input from both speech and text modalities using separate extractors tailored to the characteristics of each modality. The raw audio input is processed by an audio encoder, which extracts relevant acoustic features, while the text input is handled by the encoder of the large language model to capture linguistic features. Once these features are extracted, they are fused through various techniques, which will be elaborated upon in the following sections. After feature fusion, the combined multimodal information is sent to the large language model for final decoding and output generation. In the following sections, we will examine three distinct model architectures, each corresponding to a different input-output modality.
(a) Direct Projection
@@
(c) Combine audio and text space
@@
3.2.1 S2T
S2T models (Speech-to-Text) take speech as the input and generate text as the output. These models typically incorporate an audio encoder to extract features from the speech signal, while a text encoder is not required, as text input is absent. Therefore, to optimize model size and efficiency, a decoder-only architecture is commonly employed to generate the text output. However, aligning the audio features with the input embedding space of the large language model (decoder-only architecture) requires a conversion module. This module ensures that the audio features are properly mapped, facilitating the generation of accurate text output. As discussed in the work by Tsunoo, et al. [37], the decoder-only architecture is particularly effective for streaming end-to-end speech recognition, where efficiency and real-time performance are critical.
3.2.2 ST2S
ST2S (Speech & Text to Speech) models process both speech and text inputs by employing distinct encoders for each modality during the modal information extraction phase. Specifically, an audio encoder extracts relevant features from the speech input, while a text encoder captures features from the textual input. In the subsequent modal information fusion phase, it is essential to align the speech-derived features with the textual representation space to ensure coherent integration. The unified multimodal features are then passed to the decoder of a large language model, which carries out the inference and decoding process to generate the final output.
3.2.3 ST2ST
Speech&Text-to-Speech&Text models (ST2ST) include an additional speech modality in the output, which requires a more complex decoding mechanism during the generation phase. After modal feature extraction, modal information fusion, and large language model inference, the text output can be generated by the decoder of the large language model. However, the speech output requires a vocoder to synthesize the final speech output [38, 13].
4 Multimodal Information Fusion
Having established a preliminary understanding of the overall architecture of Speech LLMs, we will now examine the model’s key components in greater detail. Beyond the well-known stages of audio feature extraction and large language model inference, the most critical issue in Speech LLM lies in the alignment between the audio modality and the text modality. Aligning these two modalities typically requires two steps, each addressing a specific problem: The first step, which we refer to as Audio Feature Post-Process, focuses on determining what specific audio modality information is needed. The second step, called Audio and Text Connection, addresses how to effectively combine these two types of information.
4.1 Audio Feature Post-Process
In this stage, researchers currently tend to directly use the final layer output of the encoder. The current mainstream approach is to extract the output of the final layer of the encoder as the primary source of audio modality information. However, there are other various alternative methods available. For instance, some approaches involve using intermediate layer outputs to capture more granular features [39]. Others apply attention mechanisms to emphasize relevant parts of the audio signal [40].
4.2 Audio and Text Connection
Once the audio modality information is obtained, it must be integrated with the textual modality information to enable the LLM to perform the final inference. Currently, the methods used can be broadly classified into two categories: transforming the audio feature into the textual modality space, or merging the audio and textual modality spaces.
4.2.1 Audio-to-Text Modality Conversion
Converting speech modality information into the text modality space is one of the most commonly used approaches today. This method fully considers that large language models (LLMs) are primarily designed for text modalities, enabling direct projection and integration of the audio modality into the text modality space. This effectively minimizes modifications to the LLM while preserving its powerful capabilities. To achieve this, a connector or projector is needed to transform or project the extracted audio modality features. This neural network structure should minimize the loss of audio feature information while ensuring smooth transfer from the audio feature space to the text feature space.
Currently, two common methods are employed (illustrated in  Figure 4(a) and Figure 4(b)):

### v4

@@ 3.1 Speech LLMs Input-Output Modes
    The ST2T mode, the most widely used in speech understanding field, incorporates both audio and text inputs. Typically, the text input serves as an instruction or task prompt. The model processes both modalities, fusing the audio and text information to produce a final text output. This approach enables multitask capabilities and allows models to fully leverage the strengths of large language models, handling a wider variety of tasks than the S2T mode. This is also the primary mode of focus in this paper.
@@
    Currently, the ST2ST mode is widely applied in text-audio multimodal interaction scenarios through discrete sequence modeling. In simple terms, this mode requires the model to perform three key stages: speech/text understanding, text-based response generation, and speech/text output. Consequently, it demands a broader range of capabilities, such as Text-to-Speech (TTS) and Speech Translation (ST). Although ST2ST is not entirely within the scope of speech understanding, it still has significant relevance to this field. Therefore, this paper includes an analysis and comparison of this mode.
4 Model structure
Fig. 3: Overview of key stages of the Speech LLMs Workflow
@@ 4 Model structure
Various Speech LLMs architectures have been developed up to now, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference. As illustrated in Figure 3, despite differences in input and output modalities, the major stages that each model undergoes remain consistent.
All Speech LLMs begin by encoding the input information, regardless of the modality. In addition to S2T models, which take only audio as input, other models leverage pretrained audio and text encoders to obtain the representations of both modalities. The encoded audio and text embeddings are then fused to prepare them for input into the language model. This step transforms the outputs of the feature extraction process into a format compatible with the language model’s input requirements. At present, most Speech LLMs focus on generating text output; therefore, the LLM inference process typically revolves around a text decoder that receives the fused multimodal features as tokenized input to produce text. For models that optionally support audio output, such as ST2ST-type models, an LLM decoder is trained to generate audio tokens, which are subsequently converted into speech using a vocoder.
4.1 Modality Feature Extraction
As mentioned in the previous chapter, during the initial feature extraction stage, the audio is processed in two distinct ways. These approaches primarily differ in the output representation format of the audio features. Continuous Sequence Modeling, which is more commonly adopted in current Speech LLMs, encodes audio using a pre-trained and fine-tuned audio encoder, producing an audio embedding of shape [batch_size, sequence_length, feature_dimension]. In contrast, Discrete Sequence Modeling discretizes the input audio signal into a sequence of discrete audio tokens—comparable to text tokens derived from raw text—which are not represented in embedding form. In this section, we summarize common models and techniques used to extract audio features using these two approaches.
4.1.1 Continuous Sequence Modeling
Continuous sequence modeling is generally more straightforward and typically encodes audio in the form of a time-frequency representation extracted using traditional signal processing-based feature extraction methods from the raw audio signal. Given an audio input $X_{a}$, a pre-trained audio encoder is applied to generate the audio features:
$Z_{a}=g(X_{a})$ (1)
where $g$ denotes the audio encoding process. The most commonly used audio encoders in mainstream Speech LLMs are Whisper [18] and Conformer [25]. Whisper is a sequence-to-sequence Transformer model trained on a wide range of speech processing tasks, and its encoder is recognized as one of the most powerful for audio input processing. Conformer is another widely adopted architecture that combines convolutional layers with Transformers to effectively model both local and global dependencies in speech signals, particularly for automatic speech recognition tasks. Additional encoders used in mainstream Speech LLMs include WavLM [35], a predictive, self-supervised learning (SSL) pre-trained model. In this stage, aside from the audio encoder itself, subsampling modules are often applied to reduce the temporal resolution of the input features, thereby decreasing computational costs and enabling the model to focus on higher-level representations. Ultimately, the output audio feature embeddings have the shape [batch_size, sequence_length, feature_dimension], serving as the input for the next stage.
4.1.2 Discrete Sequence Modeling
Discrete sequence modeling converts raw audio into a sequence of discrete tokens that represent acoustic content and can typically be decoded back into high-quality audio. The core idea behind generating discrete tokens lies in vector quantization. Building on VQ-VAE [36], which introduced the concept of encoding continuous audio features into symbolic representations via a learned codebook, current mainstream methods for generating discrete tokens include self-supervised pretrained audio tokenizers such as HuBERT [26], and neural codec models such as EnCodec [37, 38]. A general process involves first encoding raw audio into a sequence of latent representations, as in continuous sequence modeling: $Z_{a}=g(X_{a})$, where $Z_{a}=[\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{T}]$ with $\textbf{z}_{t}\in\mathbb{R}^{d}$. Next, a quantization function $q(\cdot)$ is applied to obtain the final discrete audio tokens:
$u_{t}=q(\textbf{z}_{t})\ \text{with}\ \textbf{e}_{u_{t}}=\mathcal{Q}(\textbf{z}_{t})$ (2)
Here, $\mathcal{Q}$ denotes the quantization function that maps continuous vectors to codebook vectors $\textbf{e}_{u_{t}}\in\mathbb{R}^{d}$, resulting in $u_{t}$, the discrete token index from the vocabulary set. The final output is a sequence of discrete audio token indices: $U=[u_{1},u_{2},\dots,u_{T}]$. Specifically, self-supervised pretrained audio tokenizers, such as HuBERT, learn contextualized audio representations via masked prediction tasks and discretize intermediate features using k-means clustering. The resulting cluster indices serve as phoneme-like discrete tokens. Neural codec models, such as EnCodec, adopt an encoder–quantizer–decoder architecture in which the audio is compressed into discrete tokens using residual vector quantization and later reconstructed [39]. These models are trained end-to-end and produce high-fidelity audio that is well-suited for generative tasks. The resulting discrete audio tokens are processed alongside tokenized text, with the integration method detailed in the following section.
4.2 Modality Information Fusion
After completing the audio feature extraction stage, the most critical challenge in Speech LLMs lies in the alignment between the audio modality and the text modality. In this section, we discuss alignment methods in both continuous and discrete fashions. As introduced in the previous section, the output formats of these two methods differ, thereby requiring customized approaches to modality fusion in each case.
4.2.1 Modality Fusion of Continuous Audio Embedding
Before addressing how to effectively combine the two types of information, one important consideration is determining which specific part of the audio modality should be used. Researchers currently tend to use the output from the final layer of the encoder as the primary source of audio modality information. However, various alternative methods also exist. For example, some approaches utilize intermediate layer outputs to capture more granular features [40], while others apply attention mechanisms to emphasize relevant parts of the audio signal [35].
Once the audio embedding is obtained, it is merged into the text embedding space using two main methods:
  1. 1.
    Pre-Decoder Alignment: This category refers to methods that align audio features with text embeddings before feeding them into the LLM decoder. A common approach is to project the audio feature information into the LLM’s text feature space through a connector [33, 20]. Specifically, the tensor containing audio features typically has its own distinct feature dimension. The audio feature embedding $Z_{a}$ is projected into a vector with the same dimensionality as the text embedding using a projection function $w$: $H_{a}=w(Z_{a})$. These audio embeddings are then concatenated with the input text embedding $H_{t}$ to form a new embedding sequence that integrates both speech and text information: $H=[H_{a},H_{t}]$. This combined sequence is then fed into the LLM. Some researchers have also implicitly incorporated the projection step within the original encoder, achieving modality projection by adjusting the encoder’s parameters during training [19]. Most models use a multi-layer perceptron (MLP) as the projection module, while some adopt more advanced techniques such as Q-Former, which introduces trainable query tokens that attend to audio features and output fixed-length embeddings aligned with the LLM input space [20].
  2. 2.
    In-Decoder Alignment (Cross Attention): In addition to merging modalities before inputting them into the language model, another method establishes a bidirectional interaction between audio and text features via cross-attention [41]. In this approach, the model maintains separate audio and text streams, allowing each modality to attend to the other through cross-attention layers, thereby enabling deeper multimodal fusion within the transformer architecture.
4.2.2 Modality Fusion of Discrete Audio Tokens
For discrete audio tokens, they are typically treated in the same way as text tokens. The core challenge lies in how to process audio tokens using only a text-based vocabulary. There are essentially two ways to address this issue:
  1. 1.
    Token Mapping: The most intuitive approach is to map audio tokens to text tokens that the LLM can process. Tsunoo et al. [42] propose using CTC predictions to efficiently generate prompts—referred to as CTC prompts. Once text tokens are obtained, they are combined with tokenized input text to form a unified token sequence containing both audio and text information. This sequence is then fed into the LLM for processing. This approach not only preserves the semantic content of the audio features but also ensures consistency in the LLM’s processing pipeline.
  2. 2.
    Token Space Expansion: While projecting the speech modality into the text modality is straightforward, it does not achieve truly lossless modality fusion. Information loss and conflicts may occur during the modality conversion process (to be elaborated later). Therefore, researchers have proposed an alternative approach that modifies the input space of the large language model to natively incorporate the audio modality [30, 28, 43]. Specifically, this method augments the token space by adding audio tokens to the existing text token vocabulary, forming a unified token space. These new audio tokens are synthesized from the audio features extracted in the previous stage, thereby preserving a greater portion of the original audio information (illustrated in Figure 4(c)).
4.3 LLM Inference
The LLM inference stage consists mainly of a text decoder, which is typically an autoregressive transformer decoder that generates text tokens conditioned on input audio and/or text. Architecturally, it often mirrors standard language models like GPT, but it is specially designed or adapted to handle multimodal inputs—such as projected audio embeddings or cross-attended audio features—alongside text tokens. To accommodate modality fusion, the decoder may incorporate special tokens (e.g., <aud>, <text>) or use prompt formatting to distinguish and contextualize different modalities during generation. Despite these adaptations, the decoder follows the same autoregressive objective: predicting the next text token given previous tokens and any additional input context. In ST2ST style models, the decoder may also be extended to support the generation of discrete audio tokens. One common approach, as seen in VALL-E, treats speech generation as a conditional codec language modeling task, where the model generates audio tokens conditioned on phoneme sequences and an acoustic prompt  [44, 27]. Another approach used in SpeechGPT expands the LLM’s vocabulary to include both text and speech tokens, enabling the model to generate audio token sequences directly alongside or following text, depending on the instruction[28]. In both cases, the generated audio tokens are passed through a neural codec decoder (e.g., EnCodec[37] or HiFi-GAN[45]) to reconstruct the final waveform.
(a) Continuous Embedding Projection
@@ 4.3 LLM Inference
(c) Token Space Expansion
@@
5 Training Strategies
This section focuses on the training strategies of Speech LLMs mainly based on continuous embeddings, where speech signals are encoded into continuous feature vectors and directly consumed by a large language model. Training for these continuous-embedding Speech LLMs typically progresses through three major stages: Modality Alignment, Multitask Training, and Instruction and Preference Alignment. Each stage plays a distinct role in gradually aligning modalities, expanding task capabilities, and fine-tuning the model to follow human intent.
5.1 Modality Alignment
At the outset, Speech LLMs must learn a shared representation between spoken input and text. In this modality alignment phase, models typically leverage vast speech data to bridge the audio-text gap. A common strategy is to incorporate a pre-trained speech encoder and align its acoustic representations with a text-based LLM. For example, SALMONN and Qwen-Audio both begin with a pretrained Whisper encoder (originally trained for ASR) coupled to a LLM backbone [20, 12]. The alignment is then refined by supervised tasks that force audio and text to correspond: notably, automatic speech recognition (ASR) and audio captioning (AAC) are usually used as fundamental alignment tasks  [46, 15]. The model is trained to map an input audio sequence $a$ to a textual output $y$ (transcript or caption), effectively minimizing a cross-entropy loss
$\mathcal{L}_{\text{align}}(\theta)=-\log P_{\theta}(y\mid a),$
thereby tuning the interface between audio encoder and text decoder. In practice, only the bridging components (e.g. Q-Former modules or adapter layers) may be updated at this stage (the large language model and encoder often remain either frozen or are adapted with low-rank updates), ensuring the rich pre-trained knowledge in each modality is preserved while their features become synchronized. Self-supervised learning can further enhance alignment: some models (e.g. SEED-ASR) first pre-train the speech encoder on millions of hours of unlabeled audio to learn latent acoustic units, and then align these with text by transcribing massive paired data. Such pre-training may involve predicting masked audio segments or contrastively matching audio embeddings to pseudo-text units, building a modality-agnostic representation. The outcome of modality alignment is a foundation where spoken content is represented in the LLM’s latent space, capturing linguistic information from speech signals. This is evidenced by improved transcription and audio comprehension early in training—OSUM, for instance, reports that concurrently training an ASR task alongside any new audio task (an “ASR+X” strategy) accelerates alignment, as the model always learns to extract and decode the spoken words before handling additional subtasks. Overall, this phase imbues the model with a basic “speech understanding” ability, aligning phonetic and lexical content of audio with text tokens, which is crucial before tackling more complex or diverse tasks.
5.2 Multitask Training
With modalities aligned, the next phase develops broad audio-language capabilities via multitask training. Here, a single model is jointly trained on a wide variety of supervised tasks, spanning speech transcription, translation, and higher-level audio understanding. The training data consist of many triples $(a,p,y)$: an audio input $a$, a textual prompt or task descriptor $p$, and the expected textual output $y$. The model is optimized to predict $y$ from $(a,p)$ for all tasks in a unified fashion. Formally, if $\mathcal{T}$ is the set of task types, one can define a combined objective as the weighted sum of task-specific losses:
$\mathcal{L}_{\text{multi}}(\theta)=\sum_{t\in\mathcal{T}}\lambda_{t}\;\mathbb{E}_{(a,p,y)\sim D_{t}}\big{[}-\log P_{\theta}(y\mid a,p)\big{]},$
where $D_{t}$ is the training set for task $t$ and $\lambda_{t}$ balances the tasks. In practice, many Speech LLMs implement this by converting each task into an instruction-following format. For example, Qwen-Audio introduces special tokens or natural language prompts that indicate the task (<Transcribe>, <Translate>, “Describe the sound:”, etc.), so that the LLM can differentiate and perform over 30 diverse audio tasks. By co-training on multilingual ASR, speech-to-text translation, speaker identification, emotion classification, audio question-answering, music tagging, and more, the model acquires multi-faceted knowledge. This joint learning promotes generalization: shared acoustic-linguistic features are reused across tasks, and training signals from simpler tasks (e.g. phonetic recognition) assist more complex ones (e.g. spoken Q&A). In Qwen-Audio’s case, hierarchical task tags were used to encourage knowledge sharing while avoiding label conflicts, and the resulting model achieved state-of-the-art results on many benchmarks without task-specific fine-tuning. Academic models like OSUM followed a similar path with fewer resources: OSUM combined eight speech understanding tasks in training (from ASR to speaker attributes), always pairing each target task with an ASR objective. This ensured that the model never loses grounding in literal transcription, yielding more stable convergence and better overall performance. SALMONN likewise performed an instruction-based multitask fine-tuning (after initial alignment) on a wide range of speech, audio, and music understanding tasks. Notably, SALMONN’s multitask stage endowed it with numerous skills (recognizing speech, translating it, answering questions about audio, classifying emotions, etc.), effectively turning the LLM into a general audio assistant. A strength of this phase is the emergence of cross-task capabilities: since the model sees varied audio-text pairs, it can handle composite instructions and scenarios that were not explicitly in training. For instance, a multitask-trained model might learn to perform spoken language identification implicitly while doing translation, or combine speech recognition and reasoning to summarize an audio clip. By the end of multitask training, the Speech LLM has become a versatile audio-informed model with competencies across many domains, all encoded within one unified network.
5.3 Instruction and Preference Alignment
The final development phase focuses on aligning the model’s behavior with user instructions and preferences, ensuring that it follows natural language commands reliably and produces responses that are helpful and context-appropriate. This stage often begins with supervised instruction tuning. The model is fine-tuned on curated instruction-response datasets, where each example is a conversation or prompt $X$ (which may include an audio input and textual query) paired with a high-quality response $Y$. The fine-tuning objective is to minimize the negative log-likelihood of the reference response given the prompt:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(X,Y)\sim D_{\text{instr}}}\left[\sum_{t=1}^{|Y|}\log P_{\theta}(y_{t}\mid y_{<t},X)\right].$
This supervised fine-tuning (SFT) teaches the model to produce coherent, instruction-following answers in a conversational style. For example, Qwen2-Audio undergoes an SFT stage to become an interactive assistant (“Qwen2-Audio-Chat”), trained on multi-turn dialogues that involve both voice and text instructions. The effect is to instill the model with the format of polite, helpful responses, and the ability to handle follow-up questions or mixed-modality inputs (such as a user speaking a query in audio followed by a text instruction). SALMONN similarly leveraged instruction tuning on assorted audio-text tasks formatted as user prompts, which improved its adherence to queries; however, the authors found that this alone caused the model to sometimes over-fit to the tasks seen (producing, say, verbatim transcriptions even for novel instructions). To mitigate this, SALMONN introduced a brief activation tuning step: a few-shot fine-tuning on prompts requiring more creative, open-ended outputs (like audio-based storytelling), which helped restore the model’s generalization and “emergent” abilities beyond the narrow supervised tasks.
After supervised alignment, some advanced Speech LLMs apply preference alignment using reinforcement learning  [12]. This step optimizes the model’s responses according to human feedback or a learned reward function, refining qualities such as factual accuracy, relevance, or user satisfaction [47, 48]. A prominent method is reinforcement learning from human feedback (RLHF) operationalized via Proximal Policy Optimization (PPO) [49]. In RLHF, the model is treated as a policy $\pi_{\theta}(y|X)$ that generates a response $y$ to an input $X$, and a reward model $R(X,y)$ (trained from human preference rankings) scores the response. The goal is to maximize expected reward $E_{X}[E_{y\sim\pi_{\theta}}[R(X,y)]]$ while keeping the language model’s output distribution close to the pre-trained model to avoid degrading fluency. PPO achieves this by updating $\theta$ in the direction of the policy gradient $\nabla_{\theta}J\approx E_{X,y}[R(X,y)\,\nabla_{\theta}\log\pi_{\theta}(y|X)]$ with constraints on the KL-divergence from the original policy. In practice, the PPO update uses a clipped surrogate objective to ensure stable improvements. An alternative, lightweight approach is Direct Preference Optimization (DPO)  [50], which was employed in Qwen2-Audio’s final training. DPO sidesteps training a separate reward network by directly using preference pairs $(y^{+},y^{-})$ for a given prompt $X$. It adjusts the model to increase the probability of preferred response $y^{+}$ over the disfavored $y^{-}$. One can frame the DPO loss as:
$\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(X,y^{+},y^{-})}\left[\log\sigma\!\Big{(}\frac{1}{\beta}(\log\pi_{\theta}(y^{+}|X)-\log\pi_{\theta}(y^{-}|X))\Big{)}\right]\!,$
where $\sigma$ is the sigmoid function and $\beta$ a temperature hyperparameter. Minimizing this loss pushes $\pi_{\theta}$ to assign higher probability to the superior response $y^{+}$, aligning the model with human preferences without the complexity of a reward model loop. By the end of this reinforcement or preference alignment phase, the Speech LLM is highly tuned to follow instructions in a human-preferred way: it will listen to a user’s voice request and respond with appropriate content, tone, and detail. Crucially, this stage often improves the model’s factuality and safety (by penalizing incorrect or undesirable continuations) – Qwen2-Audio’s authors note that after DPO fine-tuning, the model more faithfully adheres to user intent and exhibits fewer mistakes in complex audio-centric instructions. The strength of the instruction alignment phase lies in making the model not just capable, but also reliably usable: it transforms the multitask-trained foundation into a polite conversational agent that can handle open-ended voice interactions. However, it is worth noting that this phase has received limited attention in current Speech LLM research focused on speech understanding. Most existing works primarily emphasize modality alignment and task-specific performance, while instruction-following behavior, dialogue consistency, and safety alignment remain underexplored. As Speech LLMs continue to evolve toward general-purpose audio agents, further investigation into instruction and preference alignment is expected to play a crucial role in improving usability, controllability, and user interaction quality.
All of these stages combined – modality alignment, broad multitask learning, and final alignment with instructions/preferences – produce state-of-the-art Speech LLMs that can understand speech in rich context and interact with users naturally, bridging the gap between spoken language and the powerful reasoning of text-based LLMs.
(a) Combine audio and text space
(b) Combine audio and text space
Fig. 5: Two ways of generating SFT data for Speech LLMs training
6 Datasests
Speech Large Language Models (Speech LLMs) are increasingly benefiting from the expanding variety of datasets. In this section, we categorize and analyze these datasets from two perspectives: the source of the data and its structural composition.
6.1 Datasets Based on Real-Recorded Audio
6.1.1 Manually Annotated Datasets
Traditional Single-Task Datasets
A significant portion of the datasets used to train Speech LLMs consists of audio-text pairs. Numerous traditional single-task datasets are readily available, providing audio-text pairs for training. The most common manually annotated datasets that developed for traditional Speech tasks, include Automatic Speech Recognition (ASR), Speech Translation(ST), Speaker Diarization (SD), and other single-task-oriented datasets. These datasets are widely used across various branches of speech understanding due to their early development and large availability. Besides using these tasks to pre-train LLM, these datasets can also be directly used to train downstream tasks for finetuning the Large Language Models.
The most commonly used datasets are as follows: